I was hoping that you could help me solve this quadratic problem using the quadratic formula.
2x^2 - 5x = 3
Any help on this problem would be helpful to me as I am not understanding these problems very well. Thank you in advance for your help.

Answer by stanbon(75887)

2x^2 - 5x = 3
-----------------
2x^2-5x-3 = 0
x = [5 +- sqrt(25-4*2*-3)]/4
x = [5 +- sqrt(49)]/4
x = [ 5 +- 7]/4
x = 3 or x = -1/2
========================
Cheers,
Stan H.